OpenStudy (anonymous):

Need help with functions! one of the problems is asking is it is a function or not: f(x)=2x+10 and also f(x)=5 ... how do i solve this? Also, I have to graph f(x)=x^2-4 ... there are no given numbers for x so how do i solve this and graph it????

OpenStudy (anonymous):

it is a function f(x)=y=2x+10 f(x)=y=5 -> 2x+10=5 -> x= (5-10)/2 y= x^2-4... the basic solution is you substitute any x ( for example 1,2,3..) into the equation to get for y. Then you can graph it yourself

OpenStudy (anonymous):

Thank you